OpenStudy (megannicole016):

please help me im so confused with Algebra :( An expression is shown below: f(x)=-16x^2+8x+3 Part A: What are the x-intercepts if the graph of the f(x)? Show your work. Part B: Is the vertex of the graph of f(x) going to be a maximum or minimum? What are the coordinates of the vertex? Justify your answers and show your work. Part C: What are the steps you would use the graph f(x)? Justify that you can use the answers obtained in Part A and Part B to draw the graph.

OpenStudy (blurbendy):

you can simplify this function to be f(x) = -(4x-3)(4x+1) 4x - 3 = 0 x = 3/4 4x +1 = 0 x = -1/4 Can you figure out B?